OpenStudy (anonymous):

Find the product of (x + 3)^2. x^2 + 6x + 9 x^2 − 6x + 9 x^2 − 9 x^2 + 9

OpenStudy (zehanz):

Write as (x+3)(x+3) and then FOIL

OpenStudy (hba):

Use(a+b)^2=a^2+2ab+b^2

OpenStudy (anonymous):

so D?

OpenStudy (zehanz):

No, you are missing out something here: (x+3)(x+3)=x²+3x+3x+3²=...

OpenStudy (zehanz):

It is even quicker to use @hba's tip: (a+b)²=a²+2ab+b². Just set a=x and b=3.

OpenStudy (anonymous):

SO A would be right then

OpenStudy (zehanz):

Yes! Promise me you'll learn (a+b)²=a²+2ab+b², because it really helps finding answers quickly ;)

OpenStudy (anonymous):

yes i will thank you. its us the signs that confuse me cus people tell me u reverse the sign and all kinds of stuff lol

OpenStudy (zehanz):

yw!